# Modeling the TMD shape function in $J/\psi$ electroproduction

### v2

@@ 1 Introduction
Despite decades of progress, many physical aspects of quarkonium dynamics remain poorly understood; for comprehensive discussions we refer to recent reviews Boer:2024ylx ; Lansberg:2019adr ; Andronic:2015wma . A central open question concerns the mechanism governing the production and decay of heavy-quark bound states. Several NP models and effective field theories have been proposed, among which non-relativistic QCD (NRQCD) Bodwin:1994jh has emerged as the most extensively studied framework. NRQCD, as an effective field theory of QCD, provides a systematic factorization of quarkonium cross-sections into two distinct components: short-distance coefficients and long-distance matrix elements (LDMEs) Nayak:2005rt ; Nayak:2005rw ; Kang:2011zza ; Kang:2014tta . The former encodes the perturbatively calculable production of a heavy-quark pair at short distance. In contrast, the LDMEs describe the NP transition of the heavy-quark pair into a physical quarkonium state, characterized by its spin $S$, orbital angular momentum $L$, total angular momentum $J$, and color configuration $[c]$ (octet $[8]$ or singlet $[1]$), conventionally denoted as $n=\,^{2S+1}L_{J}^{[c]}$. Since LDMEs encapsulate hadronization probabilities, they cannot be computed perturbatively and must instead be extracted from experimental data Butenschoen:2011yh ; Shao:2014yta ; Bodwin:2015iua ; Feng:2018ukp ; Brambilla:2022ayc . Furthermore, not all heavy-quark configurations contribute equally to quarkonium formation, as the LDMEs scale with powers of the heavy-quark relative velocity v Braaten:1996ix , such that practical calculations are truncated at finite order in the v-expansion. Nonetheless, theoretical uncertainties persist, most notably the apparent process dependence of the extracted LDMEs, which challenges their assumed universality within NRQCD.
@@
While a detailed discussion of these issues lies beyond the scope of the present work, they nonetheless provide important context for the study of quarkonium production at low transverse momentum that we pursue here. In this case, a modification of NRQCD becomes necessary, specifically the investigation of the role played by soft gluons within the theoretical framework. To address this, two approaches have been developed: pNRQCD Brambilla:1999xf ; Brambilla:2004jw and vNRQCD Luke:1999kz ; Rothstein:2018dzq . In this work we focus on vNRQCD, where two types of gluons, ultrasoft and soft, are explicitly present. Within this theoretical framework, a power counting of the interactions among the relevant degrees of freedom is carried out, such that heavy quarks can only undergo off-shell fluctuations at the ultrasoft scale, implying that single soft-gluon emissions from a heavy-quark line do not occur at leading power. Therefore, soft gluons are not treated as dynamical degrees of freedom, but only as mediators whose self-interactions are already encoded in the construction of the effective field theory.
@@
Recent developments Echevarria:2019ynx ; Fleming:2019pzj ; Echevarria:2024idp ; Copeland:2025vop have clarified that a consistent formulation of TMD factorization in quarkonium production requires incorporating the interplay between soft radiation and bound-state formation, the latter being encoded in a process-dependent TMD shape function (TMDShF) or in a recently-defined TMD soft transition function (TMDSTF). The necessity of introducing TMDShF was also demonstrated through a matching analysis of the $J/\psi$ electroproduction cross-section between the intermediate- and low-transverse-momentum regimes Boer:2020bbd ; Boer:2023zit . In that framework, an explicit expression for the NLO TMDShF was obtained; however, it does not coincide with the definition derived via the operator-level construction in Ref. Echevarria:2024idp . The origin of this discrepancy can be traced to a different choice of hard scale, a point that will be revisited in the current text. Regarding the TMDSTF, while this work was being finalized, Ref. Copeland:2025vop appeared, introducing the so-called TMDSTF. This object encodes the transition of color-octet (CO) heavy-quark pairs into color-singlet configurations through the emission of soft gluons. In the power counting of vNRQCD, the operator defining the TMDSTF is expected to be leading with respect to the TMDShFs, since the color-singlet LDME scales as v3, whereas the color-octet LDMEs scale as v7. At the same time, the presence of the gluon field renders the operator subleading in the strong coupling. However, in this case, the strong coupling must be evaluated at the soft scale which is typically larger than its value at the hard scattering scale. These considerations highlight the need for a more systematic investigation of the interplay between the relevant power-counting parameters, as well as of the possible dominance of TMDSTF contributions over those from TMDShFs, or viceversa, in the cross-section. At present, the TMDSTF has only been computed perturbatively at leading order and without including TMD evolution, preventing any firm conclusions. Our preliminary expectation, based on the numerical results shown in Ref. Copeland:2025vop , is that the TMDSTF contribution to the cross-section is strongly suppressed relative to the $^{1}S_{0}^{[8]}$ TMDShF contribution, owing to its behavior at small values of $b_{T}$. Nevertheless, as emphasized in Ref. Copeland:2025vop , this behavior is expected to be significantly modified once TMD evolution is consistently implemented.
@@
The central aim of the present work is to provide predictions for the unpolarized $J/\psi$ electroproduction cross-section at low transverse momentum at the future Electron-Ion Collider AbdulKhalek:2021gbh . To this end, we adopt the TMD factorization framework developed in Ref.  Echevarria:2024idp , in which the cross-section is expressed as the convolution of a TMD gluon distribution with the operator-defined TMDShF. Since the next-to-leading order (NLO) hard function has not previously been computed within this framework for the process under study, we also perform this calculation. Concretely, we analyze electroproduction in the small-transverse-momentum region at lowest perturbative order, $\mathcal{O}(\alpha_{em}\alpha_{s})$. This corresponds to the color-octet states $^{1}S_{0}^{[8]}$, $^{3}S_{1}^{[8]}$, and $^{3}P_{J}^{[8]}$ (see, e.g., Ref. Schuler:1997is ), collectively denoted as the $\mathcal{O}(\alpha_{s})$ CO channel.
@@ 2 Cross-section in TMD factorization
where we have summed over the spin of the final lepton, and $\lambda_{\ell}=\pm 1$ is twice the helicity of the incoming lepton.
@@
where $J^{\mu}$ is the electromagnetic current in QCD. After the TMD factorization procedure, which is detailed in Ref. Echevarria:2024idp , we obtain that the hadron tensor, in transverse-momentum space, can be factorized in terms of the [n] hard function $H_{[n]}$, [^1] In the current work, we denote the configuration of the heavy-quark pair as $[n]$ instead of $n$, avoiding confusion with the $n$-light-collinear direction. the gluon TMDPDF $G_{g/N}$ and the $[n]$-TMDShF $S_{[n]\to J/\psi}$:
@@
where $\mu_{H}$ is the hard factorization scale, $\zeta$ is the rapidity scale, such that $\zeta_{A}\zeta_{B}=\mu_{H}^{2}$, and $\mu$ is renormalization scale. The tensor $\Gamma^{\nu\alpha}_{[n]}$ refers to the tensorial structure of the partonic process through the heavy-quark pair is produced; as can be seen, it depends on the state in which the pair is formed and is common to all orders in $\alpha_{s}$.
@@ 2.1 NLO hard scattering for electroproduction
As no explicit next-to-leading order (NLO) calculation of the virtual contribution to the electroproduction cross-section is available in the literature, we have performed this calculation ourselves and present the results here. Our approach follows the procedure of Ref. Maltoni:1997pt , where the corresponding process was evaluated for photoproduction. In our case, the introduction of an additional scale, $Q^{2}$, renders the calculation more intricate; for this reason, the details are provided in Appendix B.
@@
is a combination of the two scales involved in the hard scattering of leptoproduction. [^2] We note this function does not coincide with the hard scale adopted in previous studies of quarkonium electroproduction, where the virtual corrections to the cross-section at NLO were not computed explicitly. In particular, Ref. Boer:2023zit illustrates this ambiguity: there, the TMD shape function was extracted via a matching procedure between the low- and intermediate-transverse-momentum regimes, yielding a distribution that differs from the operator-level definition of the TMDShF given in Ref. Echevarria:2024idp . The origin of this discrepancy lies in the choice of hard scale: Ref. Boer:2023zit employs $\sqrt{M^{2}+Q^{2}}$ instead of the expression in Eq. (10), leading to a TMDShF with an unphysical $Q^{2}$ dependence. This in turn motivated a dispensable discussion regarding how to remove such dependence, noting that the TMDShF is intrinsically a soft distribution. We have explicitly verified that adopting $\mu_{H}$ as in (8) in the analysis of Ref. Boer:2023zit eliminates the spurious $Q^{2}$ dependence, yielding a TMDShF that agrees with the operator-level result of Ref. Echevarria:2024idp . Here the strong coupling is bare, that is, the UV divergence disappears after renormalization. Since the IR divergences must match in both the effective theory and QCD, it can be easily verified that these divergences arise purely from the contribution of the gluon TMDPDF and the TMDShF at NLO. The remainder will be a finite term that contains no divergences and constitutes the hard function:
@@
Here the coefficients $f_{[n]}$ are given by rational functions of $\rho\equiv Q^{2}/M^{2}$ and depend on the Casimir operators $C_{A}$ and $C_{F}$, while $I_{1}$ and $I_{2}$ denote one-loop integrals that we have evaluated numerically; their explicit forms are provided in Appendix B. The corresponding expression for $H^{(1)}_{[n]}$ in photoproduction is readily obtained by taking the limit $Q\to 0$ (i.e., $\rho\to 0$) of the above result, which reproduces the known result in the literature, see e.g. Ref. Maltoni:1997pt . [^3] One should notice that the hard function corresponding to the $^{3}P_{1}^{[8]}$ state develops a singularity of the form $1/(M^{2}-Q^{2})$ at $M^{2}=Q^{2}$. Since our analysis focuses on the kinematic region where $z\to 1$, the condition $Q^{2}>M^{2}$ appears physically sensible, as for $Q^{2}<M^{2}$ the available energy would be insufficient to produce the final state. Given that our results for the $^{3}P_{0}^{[8]}$ and $^{3}P_{2}^{[8]}$ channels are consistent with the literature in the photoproduction limit, we restrict our cross-section analysis and subsequent EIC predictions to the region $Q^{2}\gtrsim M^{2}$.
@@ 3.1 Evolution and non-perturbative behavior
Here $\left<\mathcal{O}^{[m]}\right$> denotes the $[m]$-LDME, $C_{[m]}^{[n]}$ is the matching coefficient relating the $[n]$-TMDShF to the $[m]$-LDME (see Appendix A), and $N_{\text{pol}}^{(J)}=2J+1$ is the number of polarizations of the state $[m]$ with total angular momentum $J$. The renormalization scale $\mu$ and the rapidity scale $\zeta_{B}$ are defined at the point where the non-perturbative input for the TMDShF is specified.
@@
Within the TMD evolution factor $R$, $\gamma_{S}$ is called the TMDShF anomalous dimension and $\mathcal{D}$ is called also the rapidity anomalous dimension. The latter describes the soft-gluon exchange between hadrons and can be determined from the rapidity divergent part of the soft function at small values of $b_{T}$ Echevarria:2015byo . In fact, its perturbative behavior, named $\mathcal{D}_{pert}$, is currently known at N3LO Duhr:2022yyp . We emphasize that LDMEs must evolve from the scale at which they are extracted to the scale $\mu_{i}$, as we discuss below. Moreover, since the anomalous dimensions of the two scales are correlated, the mutual dependence is explicitly worked out through the so-called cusp anomalous dimension $\Gamma_{cusp}$, such as the TMDShF anomalous dimension can be written as
@@
where $\Gamma_{cusp}$ is currently known at N4LO Herzog:2018kwj . The anomalous dimension $\gamma_{s}$ refers to the finite part of the renormalization of the shape function and is known at NLO Echevarria:2024idp . We direct the reader to the Appendix A, where there is a summary of all these expressions in the truncated perturbation theory.
@@
where $b_{c}$ fixes the lower limit $b_{T\text{min}}\equiv b_{0}/\mu_{H}\leq b_{T}$ which marks the point where $\mu_{b}\equiv b_{0}/b_{T}$ becomes larger than $\mu_{H}$. Therefore, this $b_{T}$-prescription ensures that $b_{T\text{min}}\leq b_{T}\leq b_{T\text{max}}$. It is easy to check that $b_{T}^{*}$ approaches $b_{T\text{min}}$ and $b_{T\text{max}}$ slowly, and there is not a clean separation of both perturbative and NP regions, so this can affect the functions of $b_{T}^{*}$ as the shape function, such that the NP distribution contributes to the perturbative one. However, since the point where perturbation theory starts to fail is defined in the range of 0.51.50.5-1.5 GeV-1, but not a fixed value, we will not dwell on it at this point and we will use the expressions of Eq. (14).
@@
In Fig. 1, we display the evolution of the $^{1}S_{0}^{[8]}$ LDME as a function of $b_{T}$, where the evolution scale is chosen as $\mu=\mu_{b^{*}}$. The panels correspond to two representative hard scales, and the colored curved denote different input values do the P-states LDMEs entering the evolution equation of the $^{1}S_{0}^{[8]}$ state. Note that for the most of the LDME sets, the value of the P-states LDMEs are less than the S-wave LDMEs. For this analysis, we consider the SV set Sharma_2013 , where $\left<{}^{1}S_{0}^{[8]}\right$> is equal to 0.018 GeV3; although this value has been extracted at the scale $m_{T}$, here we consider the approximation $m_{T}\approx M$. The strength of this suppression is controlled by the choice of the P-wave matrix elements: small values produce only mild deviations, while larger values result in a more pronounced decrease. This sensitivity illustrates the interplay between S-wave and P-wave channels in the evolution kernel. The comparison between two panels highlights the scale dependence. The lower plots quantify the deviation with respect to the extracted $^{1}S_{0}^{[8]}$ LDME. In the intermediate and NP domains, the deviation grows, reaching up to $\mathcal{O}(10\%-20\%)$, depending on the chosen P-wave inputs. Furthermore, we found that for P-wave LDMEs exceeding $\sim 0.1~\text{GeV}^{3}$, the $^{1}S_{0}^{[8]}$ LDME becomes negative in the intermediate and large-$b_{T}$ regions. Although it is clear we do not have any extraction for these LDMEs for $J/\psi$, after this naive analysis we conclude that the evolution is sensitive to the values of the P-states LDMEs and that this sensitivity is amplified at large $b_{T}$ and higher hard scales. Therefore, for a precise phenomenological analysis, the error induced by treating the LDMEs as constants could be significant.
@@
In Fig. 2 we show the $^{1}S_{0}^{[8]}$ TMDShF in the $q_{T}$-space at $\mu_{H}=15$ GeV, and $b_{T\text{max}}=1.5$ GeV-1 at the levels of accuracy illustrated in Tab. 1. To be consistent with the factorization theorem, we calculate the theoretical curves up to $q_{T}=\mu_{H}/2$, but in the figure is shown only the scale variation up to $q_{T}=3$ GeV. For the NP parametrization, we fix $\zeta_{B,0}=1$ (i.e., $\mathcal{D}_{\text{NP}}=0$), and we fix $B_{S}=0.5$ GeV2. The running of the strong coupling is implemented at N3LO with $\Lambda_{\text{QCD}}=0.239$ GeV-1, while the flavor number scheme is the one fixing $n_{f}=4$. The bands come from varying both $C_{b^{*}}$ and $\zeta_{B,b}$ as mentioned above, and keeping the largest variation for each point in the $q_{T}$-spectrum. The bands at NNLL + NLO and at NNL + NLO are the narrower bands, as they include a higher perturbation order for $C_{[m]}^{[n]}$. At a fixed resummation order in $S_{A}$, from comparing the green and blue bands or the red and yellow bands, we observe a reasonable reduction ($\sim 48\%$) in the theoretical uncertainty. As we can see in the right picture, the most decreasing in the uncertainty is coming from varying only $\mu_{b^{*}}$, which directly affect the NLO order of $C_{[m]}^{[n]}$ through $\hbox{ln}(\mu^{2}b_{T}^{2})$. The difference with the left picture comes from the contribution of $\mathcal{D}_{pert}$ which enters in the game when $\zeta_{B,b}\neq 1$, and the thicker bands are obtained when $\zeta_{B,b}=0.5$ as it flips the sign of the $\mathcal{D}_{pert}$ term in the $S_{A}$. It is clear that the theoretical uncertainty gets reduced as we increase the resummation accuracy by including more perturbative ingredients, especially in $C_{[m]}^{[n]}$.

### v3

@@ 1 Introduction
Despite decades of progress, many physical aspects of quarkonium dynamics remain poorly understood; for comprehensive discussions we refer to recent reviews Boer and others (2024); Lansberg (2020); Andronic and others (2016). A central open question concerns the mechanism governing the production and decay of heavy-quark bound states. Several NP models and effective field theories have been proposed, among which non-relativistic QCD (NRQCD) Bodwin et al. (1995) has emerged as the most extensively studied framework. NRQCD, as an effective field theory of QCD, provides a systematic factorization of quarkonium cross-sections into two distinct components: short-distance coefficients and long-distance matrix elements (LDMEs) Nayak et al. (2005b, a); Kang et al. (2011, 2014). The former encodes the perturbatively calculable production of a heavy-quark pair at short distance. In contrast, the LDMEs describe the NP transition of the heavy-quark pair into a physical quarkonium state, characterized by its spin $S$, orbital angular momentum $L$, total angular momentum $J$, and color configuration $[c]$ (octet $[8]$ or singlet $[1]$), conventionally denoted as $n=\,^{2S+1}L_{J}^{[c]}$. Since LDMEs encapsulate hadronization probabilities, they cannot be computed perturbatively and must instead be extracted from experimental data Butenschoen and Kniehl (2011); Shao et al. (2015); Bodwin et al. (2016); Feng et al. (2019); Brambilla et al. (2023). Furthermore, not all heavy-quark configurations contribute equally to quarkonium formation, as the LDMEs scale with powers of the heavy-quark relative velocity v Braaten (1996), such that practical calculations are truncated at finite order in the v-expansion. Nonetheless, theoretical uncertainties persist, most notably the apparent process dependence of the extracted LDMEs, which challenges their assumed universality within NRQCD.
@@
While a detailed discussion of these issues lies beyond the scope of the present work, they nonetheless provide important context for the study of quarkonium production at low transverse momentum that we pursue here. In this case, a modification of NRQCD becomes necessary, specifically the investigation of the role played by soft gluons within the theoretical framework. To address this, two approaches have been developed: pNRQCD Brambilla et al. (2000, 2005) and vNRQCD Luke et al. (2000); Rothstein et al. (2019). In this work we focus on vNRQCD, where two types of gluons, ultrasoft and soft, are explicitly present. Within this theoretical framework, a power counting of the interactions among the relevant degrees of freedom is carried out, such that heavy quarks can only undergo off-shell fluctuations at the ultrasoft scale, implying that single soft-gluon emissions from a heavy-quark line do not occur at leading power. Therefore, soft gluons are not treated as dynamical degrees of freedom, but only as mediators whose self-interactions are already encoded in the construction of the effective field theory.
@@
Recent developments Echevarria (2019); Fleming et al. (2020); Echevarria et al. (2024); Copeland et al. (2025) have clarified that a consistent formulation of TMD factorization in quarkonium production requires incorporating the interplay between soft radiation and bound-state formation, the latter being encoded in a process-dependent TMD shape function (TMDShF) or in a recently-defined TMD soft transition function (TMDSTF). The necessity of introducing TMDShF was also demonstrated through a matching analysis of the $J/\psi$ electroproduction cross-section between the intermediate- and low-transverse-momentum regimes Boer et al. (2020, 2023). In that framework, an explicit expression for the NLO TMDShF was obtained; however, it does not coincide with the definition derived via the operator-level construction in Ref. Echevarria et al. (2024). The origin of this discrepancy can be traced to a different choice of hard scale, a point that will be revisited in the current text. Regarding the TMDSTF, while this work was being finalized, Ref. Copeland et al. (2025) appeared, introducing the so-called TMDSTF. This object encodes the transition of color-octet (CO) heavy-quark pairs into color-singlet configurations through the emission of soft gluons. In the power counting of vNRQCD, the operator defining the TMDSTF is expected to be leading with respect to the TMDShFs, since the color-singlet LDME scales as v3, whereas the color-octet LDMEs scale as v7. At the same time, the presence of the gluon field renders the operator subleading in the strong coupling. However, in this case, the strong coupling must be evaluated at the soft scale which is typically larger than its value at the hard scattering scale. These considerations highlight the need for a more systematic investigation of the interplay between the relevant power-counting parameters, as well as of the possible dominance of TMDSTF contributions over those from TMDShFs, or viceversa, in the cross-section. At present, the TMDSTF has only been computed perturbatively at leading order and without including TMD evolution, preventing any firm conclusions. Our preliminary expectation, based on the numerical results shown in Ref. Copeland et al. (2025), is that the TMDSTF contribution to the cross-section is strongly suppressed relative to the $^{1}S_{0}^{[8]}$ TMDShF contribution, owing to its behavior at small values of $b_{T}$. Nevertheless, as emphasized in Ref. Copeland et al. (2025), this behavior is expected to be significantly modified once TMD evolution is consistently implemented.
@@
The central aim of the present work is to investigate the TMD shape function contribution to the unpolarized $J/\psi$ electroproduction cross-section at low transverse momentum at the future Electron-Ion Collider Abdul Khalek and others (2022). To this end, we adopt the TMD factorization framework developed in Ref.  Echevarria et al. (2024), in which the cross-section is expressed as the convolution of a TMD gluon distribution with the operator-defined TMDShF. Since the next-to-leading order (NLO) hard function has not previously been computed within this framework for the process under study, we also perform this calculation. Concretely, we analyze electroproduction in the small-transverse-momentum region at lowest perturbative order, $\mathcal{O}(\alpha_{em}\alpha_{s})$. This corresponds to the color-octet states $^{1}S_{0}^{[8]}$, $^{3}S_{1}^{[8]}$, and $^{3}P_{J}^{[8]}$ (see, e.g., Ref. Schuler (1997)), collectively denoted as the $\mathcal{O}(\alpha_{s})$ CO channel. Moreover, at $\mathcal{O}(\alpha_{s})$, the contribution from $^{3}S_{1}^{[8]}$ state vanishes due to the Lorentz and spin structure of the two-vector initial state, so this channel only appears at higher order with an additional gluon in the final state.
@@ 2 Cross-section in TMD factorization
where we have summed over the spin of the final lepton, and $\lambda_{\ell}=\pm 1$ is twice the helicity of the incoming lepton. Additionally, the contribution from the anti-symmetric (parity-violating) term will vanish for the unpolarized cross section.
@@
where $J^{\mu}$ is the electromagnetic current in QCD. After the TMD factorization procedure, which is detailed in Ref. Echevarria et al. (2024), we obtain that the hadron tensor, in transverse-momentum space, can be factorized in terms of the [n] hard function $H_{[n]}$, [^1] In the current work, we denote the configuration of the heavy-quark pair as $[n]$ instead of $n$, avoiding confusion with the $n$-light-collinear direction. the gluon TMDPDF $G_{g/N}$ and the $[n]$-TMDShF $S_{[n]\to J/\psi}$:
@@
where $\mu$ is renormalization scale, and $\zeta$’s are the rapidity scales, such that $\zeta_{A}\zeta_{B}=\mu_{H}^{2}$, with $\mu_{H}$ represents the hard factorization scale. The tensor $\Gamma^{\nu\alpha}_{[n]}$ refers to the tensorial structure of the partonic process through the heavy-quark pair is produced; as can be seen, it depends on the state in which the pair is formed and is common to all orders in $\alpha_{s}$.
@@ 2.1 NLO hard scattering for electroproduction
As no explicit next-to-leading order (NLO) calculation of the virtual contribution to the electroproduction cross-section is available in the literature, we have performed this calculation ourselves and present the results here. Our approach follows the procedure of Ref. Maltoni et al. (1998), where the corresponding process was evaluated for photoproduction. In our case, the introduction of an additional scale, $Q^{2}$, renders the calculation more intricate; for this reason, the details are provided in Appendix B.
@@
is a combination of the two scales involved in the hard scattering of leptoproduction. This scale is not an ad-hoc choice, but it emerges naturally from the perturbative calculation. [^2] We note that the hard scale of the process is clearly then $(M^{2}+Q^{2})/M$. This scale does not coincide with the hard scale adopted in previous studies of quarkonium electroproduction, where the virtual corrections to the cross-section at NLO were not computed explicitly. In particular, Ref. Boer et al. (2023) illustrates this ambiguity: there, the TMD shape function was extracted via a matching procedure between the low- and intermediate-transverse-momentum regimes, yielding a distribution that differs from the operator-level definition of the TMDShF given in Ref. Echevarria et al. (2024). The origin of this discrepancy lies in the hard scale: Ref. Boer et al. (2023) employs the ansatz $\sqrt{M^{2}+Q^{2}}$, instead of the expression obtained in this work in Eq. (10), leading to a TMDShF with an unphysical $Q^{2}$ dependence. This in turn motivated a discussion regarding how to remove such dependence, noting that the TMDShF is intrinsically a soft distribution. We have explicitly verified that adopting the hard scale $\mu$ as in (8) in the analysis of Ref. Boer et al. (2023) eliminates the spurious $Q^{2}$ dependence, yielding a TMDShF that agrees with the operator-level result of Ref. Echevarria et al. (2024). Therefore, the renormalization scale in the hard factor, $\mu$, is chosen to be proportional to $\mu_{H}$ to minimize logarithms in the perturbative calculation. Here the strong coupling is bare, that is, the UV divergence disappears after renormalization. Since the IR divergences must match in both the effective theory and QCD, it can be easily verified that these divergences arise purely from the contribution of the gluon TMDPDF and the TMDShF at NLO. The remainder will be a finite term that contains no divergences and constitutes the hard function:
@@
Here the coefficients $f_{[n]}$ are given by rational functions of $\rho\equiv Q^{2}/M^{2}$ and depend on the Casimir operators $C_{A}$ and $C_{F}$, while $I_{1}$ and $I_{2}$ denote one-loop integrals that we have evaluated numerically; their explicit forms are provided in Appendix B. The corresponding expression for $H^{(1)}_{[n]}$ in photoproduction is readily obtained by taking the limit $Q\to 0$ (i.e., $\rho\to 0$) of the above result, which reproduces the known result in the literature, see e.g. Ref. Maltoni et al. (1998). [^3] One should notice that the hard function corresponding to the $^{3}P_{1}^{[8]}$ state develops a singularity of the form $1/(M^{2}-Q^{2})$ at $M^{2}=Q^{2}$. Since our analysis focuses on the kinematic region where $z\to 1$, the condition $Q^{2}>M^{2}$ appears physically sensible, as for $Q^{2}<M^{2}$ the available energy would be insufficient to produce the final state. Given that our results for the $^{3}P_{0}^{[8]}$ and $^{3}P_{2}^{[8]}$ channels are consistent with the literature in the photoproduction limit, we restrict our cross-section analysis and subsequent EIC predictions to the region $Q^{2}\gtrsim M^{2}$.
@@ 3.1 Evolution and non-perturbative behavior
Here $\left<\mathcal{O}^{[m]}\right$> denotes the $[m]$-LDME, $C_{[m]}^{[n]}$ is the matching coefficient relating the $[n]$-TMDShF to the $[m]$-LDME (see Appendix A), and $N_{\text{pol}}^{(J)}=2J+1$ is the number of polarizations of the state $[m]$ with total angular momentum $J$. The renormalization scale $\mu$ and the rapidity scale $\zeta$ are defined at the point where the non-perturbative input for the TMDShF is specified.
@@
Within the TMD evolution factor $R$, $\gamma_{S}$ is called the TMDShF anomalous dimension and $\mathcal{D}$ is called also the rapidity anomalous dimension. The latter describes the soft-gluon exchange between hadrons and can be determined from the rapidity divergent part of the soft function at small values of $b_{T}$ Echevarria et al. (2016). In fact, its perturbative behavior, named $\mathcal{D}_{pert}$, is currently known at N3LO Duhr et al. (2022). We emphasize that LDMEs must evolve from the scale at which they are extracted to the scale $\mu_{i}$, as we discuss below. Moreover, since the anomalous dimensions of the two scales are correlated, the mutual dependence is explicitly worked out through the so-called cusp anomalous dimension $\Gamma_{cusp}$, such as the TMDShF anomalous dimension can be written as
@@
where $\Gamma_{cusp}$ is currently known at N4LO Herzog et al. (2019). The anomalous dimension $\gamma_{s}$ refers to the finite part of the renormalization of the shape function and is known at NLO Echevarria et al. (2024). We direct the reader to the Appendix A, where there is a summary of all these expressions in the truncated perturbation theory.
@@
where $b_{c}$ fixes the lower limit $b_{T\text{min}}\equiv b_{0}/\mu_{H}\leq b_{T}$ which marks the point where $\mu_{b}\equiv b_{0}/b_{T}$ becomes larger than $\mu_{H}$. Therefore, this $b_{T}$-prescription ensures that $b_{T\text{min}}\leq b_{T}\leq b_{T\text{max}}$. For $b_{0}$, see Appendix A. It is easy to check that $b_{T}^{*}$ approaches $b_{T\text{min}}$ and $b_{T\text{max}}$ slowly, and there is not a clean separation of both perturbative and NP regions, so this can affect the functions of $b_{T}^{*}$ as the shape function, such that the NP distribution contributes to the perturbative one. However, since the point where perturbation theory starts to fail is defined in the range of 0.51.50.5-1.5 GeV-1, but not a fixed value, we will not dwell on it at this point and we will use the expressions of Eq. (14).
@@
In Fig. 1, we display the evolution of the $^{1}S_{0}^{[8]}$ LDME as a function of $b_{T}$, where the evolution scale is chosen as $\mu=\mu_{b^{*}}$. The panels correspond to two representative hard scales, and the colored curved denote different input values do the P-states LDMEs entering the evolution equation of the $^{1}S_{0}^{[8]}$ state. For this analysis, we consider the SV set Sharma and Vitev (2013), where $\left<{}^{1}S_{0}^{[8]}\right$> is equal to 0.018 GeV3; although this value has been extracted at the scale $m_{T}$, here we consider the approximation $m_{T}\approx M$. We assume that the P-state LDMEs are suppressed with respect to the S-state LDMEs, and we also analyze the extreme case in which they are equal. The strength of this suppression is controlled by the choice of the P-wave matrix elements: small values produce only mild deviations, while larger values result in a more pronounced decrease. This sensitivity illustrates the interplay between S-wave and P-wave channels in the evolution kernel. The comparison between two panels highlights the scale dependence. The lower plots quantify the deviation with respect to the extracted $^{1}S_{0}^{[8]}$ LDME. In the intermediate and NP domains, the deviation grows, reaching up to $\mathcal{O}(10\%-20\%)$, depending on the chosen P-wave inputs. Furthermore, we found that for P-wave LDMEs exceeding $\sim 0.1~\text{GeV}^{3}$, the $^{1}S_{0}^{[8]}$ LDME becomes negative in the intermediate and large-$b_{T}$ regions. Although it is clear we do not have any extraction for these LDMEs for $J/\psi$, after this naive analysis we conclude that the evolution is sensitive to the values of the P-states LDMEs and that this sensitivity is amplified at large $b_{T}$ and higher hard scales. Therefore, for a precise phenomenological analysis, the error induced by treating the LDMEs as constants could be significant, but we do not include it in this work. As can be seen in Eq. (42), the functional form of the evolution for the P-states is the same as for the S-states. Therefore, if we use the same range of values for the LDMEs of the D states used in this analysis, we expect the evolution to be formally identical in structure for the P -states.
@@
In Fig. 2 we show the $^{1}S_{0}^{[8]}$ TMDShF in the $q_{T}$-space at $\mu_{H}=15$ GeV, and $b_{T\text{max}}=1.5$ GeV-1 at the levels of accuracy illustrated in Tab. 1. To be consistent with the factorization theorem, we calculate the theoretical curves up to $q_{T}=\mu_{H}/2$, but in the figure is shown only the scale variation up to $q_{T}=3$ GeV. For the NP parametrization, we fix $\zeta_{B,0}=1$ (i.e., $\mathcal{D}_{\text{NP}}=0$), and we fix $B_{S}=0.5$ GeV2. The running of the strong coupling is implemented at N3LO with $\Lambda_{\text{QCD}}=0.239$ GeV-1, while the flavor number scheme is the one fixing $n_{f}=4$. The bands come from varying both $C_{b^{*}}$ and $\zeta_{B,b}$ with a 9-point variation, and keeping the largest variation for each point in the $q_{T}$-spectrum. The bands at NNLL + NLO and at NNL + NLO are the narrower bands, as they include a higher perturbation order for $C_{[m]}^{[n]}$. At a fixed resummation order in $S_{A}$, from comparing the green and blue bands or the red and yellow bands, we observe a reasonable reduction ($\sim 48\%$) in the theoretical uncertainty. As we can see in the right picture, the most decreasing in the uncertainty is coming from varying only $\mu_{b^{*}}$, which directly affect the NLO order of $C_{[m]}^{[n]}$ through $\hbox{ln}(\mu^{2}b_{T}^{2})$. The difference with the left picture comes from the contribution of $\mathcal{D}_{pert}$ which enters in the game when $\zeta_{B,b}\neq 1$, and the thicker bands are obtained when $\zeta_{B,b}=0.5$ as it flips the sign of the $\mathcal{D}_{pert}$ term in the $S_{A}$. It is clear that the theoretical uncertainty gets reduced as we increase the resummation accuracy by including more perturbative ingredients, especially in $C_{[m]}^{[n]}$.